# Communication-Efficient Decentralized Optimization via Double-Communication Symmetric ADMM

### v1

@@ 3.3 Communication Strategy
Crucially, our design partitions the variables into two interdependent groups: $(u,w_{2})$ and $(v,w_{1})$. Information from one group is not used directly in its own update but enables the update of the other group, creating a feedback structure. This interleaving induces a coupled communication-update mechanism where each block drives progress in the other.
@@
Furthermore, this structure makes Symmetric ADMM not just suitable but essential: it provides accelerated convergence without increasing communication, and leads to a clean, symmetric algorithm. To our knowledge, this tightly coupled update-communication framework is novel in the decentralized optimization literature.
Remark 2.
Dividing the Lagrange multiplier $\lambda$ into two independent blocks $w_{1}$ and $w_{2}$ greatly enhances flexibility in communication design since the updates of two blocks are parallel to each other and realigning the order of them is allowed.
@@ 3.4 Algorithm
We denote our proposed algorithm as DS-ADMM, which encapsulates several key design features: it adopts the Symmetric ADMM framework for Decentralized optimization, incorporates a Double communication structure within each iteration, and exhibits a Symmetric structure in its use of two interleaved variable blocks.
We now present the algorithm in its fully decentralized form. For clarity, we reposition the update of the dual variable $w_{2}$—originally placed at the end of each iteration—to the beginning. The step-size parameters are set as $0<r\leq 1$ and $s=1$.
Algorithm 1 DS-ADMM
1:Initialize: $u_{i}^{(0)}=v_{i}^{(0)}=w_{1i}^{(0)}=w_{2i}^{(-\frac{1}{2})}=0$ for all $i\in\{1,\dots,n\}$ for all $i\in\{1,\dots,n\}$, mixing matrix $W\in\mathbb{R}^{n\times n}$.
@@
4:  [Communication 1 ] Transmit $a_{i}^{(t+1)}=w_{2i}^{(t+\frac{1}{2})}+\frac{1}{r}(w_{2i}^{(t+\frac{1}{2})}-w_{2i}^{(t)})$ and $u_{i}^{(t+1)}$.
@@
6:  [Communication 2] Transmit $v_{i}^{(t+1)}$ and $b_{i}^{(t+1)}=2w_{1i}^{(t+1)}-w_{1i}^{(t+\frac{1}{2})}$.
@@ 4 Convergence Analysis
which is positive definite. We also define the concatenated variable $w=(u^{\top},v^{\top},\lambda^{\top})^{\top}\in\mathbb{R}^{4nd}$.
@@ 4.1 General Sublinear Convergence
Theorems 3.3 and 4.2 of (PRSM, ) imply that the proposed algorithm enjoys a general sublinear convergence rate $\mathcal{O}(1/t)$ without requiring strong assumptions on the objective functions.
@@
Let $\{w^{(t)}\}$ be the sequence generated by DS-ADMM. Then $\{w^{(t)}\}$ converges to a solution point $w^{\infty}$, and the following non-ergodic sublinear rate holds:
$\|w^{(t)}-w^{(t+1)}\|^{2}\leq\frac{1}{\beta\tau(t+1)}\cdot\frac{1+r}{1-r}\left(\|w^{(1)}-w^{(0)}\|_{H}^{2}+\|v^{(1)}-v^{(0)}\|_{Q}^{2}\right).$ (5)
@@ 4.2 Linear Convergence under Metric Subregularity
Although various results on linear convergence of ADMM and its variants exist (e.g., (PADMM, ; ADMMconv1, ; ADMMconv, ; GSADMMConv, ; PRSM, )), none directly apply to our algorithm. Nevertheless, we adapt ideas from these works to establish linear convergence under a standard regularity condition known as metric subregularity.
@@
$T_{\text{KKT}}(w):=\begin{pmatrix}\partial f(u)-A^{\top}\lambda\\ \partial g(v)+B^{\top}\lambda\\ Au-Bv\end{pmatrix}$, (7)
and solution set $\Omega^{*}:=\{w\mid 0\in T_{\text{KKT}}(w)\}$.
@@
Suppose $T_{\mathrm{KKT}}$ is metrically subregular at $(\bar{w},0)$ with modulus $c$ for any $\bar{w}\in\Omega^{*}$. Then the sequence $\{w^{(t)}\}$ generated by DS-ADMM converges Q-linearly to $\Omega^{*}$, i.e., there exist integer $T>0$ and constant $\epsilon>0$ such that for all $t>T$:
$\mathrm{dist}^{2}_{H}(w^{(t+1)},\Omega^{*})\leq\frac{1}{1+\epsilon}\cdot\mathrm{dist}^{2}_{H}(w^{(t)},\Omega^{*}),$ (8)
@@
$\epsilon=\frac{\phi}{c^{2}\delta\theta}>0,\quad\phi=\min\left\{2\beta\rho,\frac{1-r}{\beta}\right\},$ (9)
@@
$\delta=\max\left\{6r^{2}+\frac{2}{\beta^{2}},\,12\beta^{2}+4+(\tau\beta)^{2},\,3(\tau\beta)^{2}\right\}\quad\theta=\frac{2r^{2}\beta^{2}+1}{\beta(r+1)}+(2+\tau-r)\beta.$ (10)
@@ 4.3 Sufficient Conditions for Metric Subregularity
The following proposition gives a characterization of the metric subregularity of $T_{\mathrm{KKT}}$. Each case is justified by different results from the literature: Theorem 46 and Theorem 60 of (ADMMconv, ) support the first condition, Robinson’s continuity property (Robinson, ) establishes the second, and Lemma 4 of (PD, ) together with Theorem 60 of (ADMMconv, ) imply the third.
@@
The KKT mapping $T_{\mathrm{KKT}}$ is metrically subregular at $(\bar{w},0)$ for any $\bar{w}\in\Omega^{*}$ under any of the following conditions: (i) each $f_{i}$ is smooth and strongly convex, and each $g_{i}$ is PLQ; (ii) all $f_{i}$ and $g_{i}$ are PLQ; (iii) all $f_{i}$ and $g_{i}$ are smooth and strongly convex.
@@ 5 Numerical Experiments
We evaluate the performance of our proposed algorithm on standard decentralized composite optimization problems with smooth–nonsmooth structure, where either the objective or penalty function is nonsmooth. Comparisons are made against four representative methods: Decentralized Proximal ADMM (DPADMM, ), PG-EXTRA (PG-EXTRA, ), NIDS  (NIDS, ) and ProxMudag (ProxMudag, ). All experiments are conducted on a machine equipped with an Intel Core i7-1260P CPU and 16GB RAM.
Throughout the experiments, the number of agents is fixed at $n=30$. The communication network is modeled as a random graph with edge probability $p=0.5$, and the corresponding mixing matrix $W$ is constructed using Metropolis-Hastings weights (see Appendix A for details). Each dataset is partitioned evenly among the agents to define local loss functions. The global regularization term $g(x)$ is split uniformly as $g_{i}(x)=\frac{1}{n}g(x)$ for all agents.
All adaptive parameters (such as step sizes and penalty coefficients) are tuned to ensure the best empirical performance. Non-adaptive parameters follow the choices and guidelines suggested in the original works of the respective baseline methods. For our algorithm, we use fixed dual step sizes $(r,s)=(0.99,1)$ and a proximal coefficient $\tau=0.01$.
We measure suboptimality using $F(\bar{u}^{(t)})-F(u^{*})$, where $\bar{u}^{(t)}$ is the average of local primal variables at iteration $t$, and $u^{*}$ is a centralized optimal solution computed to high precision.
5.1 Lasso Regression
Figure 1: Performance on the Lasso regression task. Left: suboptimality vs. iterations. Right: suboptimality vs. communication rounds.
We first test the algorithms on a Lasso regression task, which consists of a quadratic loss and an $\ell_{1}$-regularization term. We use the a9a dataset from the LIBSVM repository (Libsvm, ). The objective function is given by:
$f_{i}(x)=\frac{1}{2m}\|A_{i}x-b_{i}\|^{2},\quad g_{i}(x)=\frac{\lambda}{n}\|x\|_{1},$
where $A_{i}$ and $b_{i}$ represent the local data on agent $i$, and $m$ is the total number of training samples. Here, $f_{i}$ is smooth and strongly convex, while $g_{i}$ is convex but nonsmooth. The regularization parameter is set to $\lambda=\frac{1}{m}$.
Figure 1 compares performance in terms of both iteration count and total communication rounds. Our proposed algorithm demonstrates significantly lower computational cost and reduced communication overhead compared to existing methods on this task.
5.2 SVM Classification
Next, we evaluate the algorithms on an $\ell_{2}$-regularized SVM classification problem, which involves hinge loss and an $\ell_{2}$-norm regularizer. The a1a dataset from LIBSVM is used. The objective function is:
$f_{i}(x)=\frac{1}{m}\sum_{j\in S_{i}}\max(0,1-b_{j}a_{j}^{\top}x),\quad g_{i}(x)=\frac{\lambda}{2n}\|x\|_{2}^{2},$
where $S_{i}$ is the local data index set on agent $i$, and $m$ is the total number of samples. In this setting, $f_{i}$ is convex but nonsmooth, and $g_{i}$ is smooth and strongly convex. We again use $\lambda=\frac{1}{m}$.
Note that ProxMudag is not included in this comparison because it requires the nonsmooth term to be globally coupled, which is incompatible with this separable formulation (ProxMudag, ). The results in Figure 2 show that our method achieves both faster convergence and lower communication costs.
Figure 2: Performance on the SVM classification task. Left: suboptimality vs. iterations. Right: suboptimality vs. communication rounds.
Additional experimental results of graph with $p=0.2$ are provided in Appendix D. These results demonstrate that DS-ADMM consistently outperforms existing methods across both tasks and further validate the linear convergence guarantees presented in our theoretical analysis.
@@ Appendix B Global Form of Update
Then we decompose the Lagrange multiplier $\lambda$:
@@
$\left\{\begin{aligned} u^{(t+1)}&=\arg\min_{u\in\mathbb{R}^{nd}}f(u)-\langle w_{1}^{(t)},\tilde{W}u\rangle-\langle w_{2}^{(t)},u\rangle+\frac{\beta}{2}\|\tilde{W}u-v^{(t)}\|^{2}+\frac{\beta}{2}\|u-\tilde{W}v^{(t)}\|^{2}\\ &+\frac{1}{2}\|u-u^{(t)}\|_{Q}^{2},\\ w_{1}^{(t+\frac{1}{2})}&=w_{1}^{(t)}-r\beta(\tilde{W}u^{(t+1)}-v^{(t)}),\quad w_{2}^{(t+\frac{1}{2})}=w_{2}^{(t)}-r\beta(u^{(t+1)}-\tilde{W}v^{(t)}),\\ v^{(t+1)}&=\arg\min_{v\in\mathbb{R}^{nd}}g(v)+\langle w_{1}^{(t+\frac{1}{2})},v\rangle+\langle w_{2}^{(t+\frac{1}{2})},\tilde{W}v\rangle+\frac{\beta}{2}\|\tilde{W}u^{(t+1)}-v\|^{2}\\ &+\frac{\beta}{2}\|u^{(t+1)}-\tilde{W}v\|^{2}+\frac{1}{2}\|v-v^{(t)}\|_{Q}^{2},\\ w_{1}^{(t+1)}&=w_{1}^{(t+\frac{1}{2})}-s\beta(\tilde{W}u^{(t+1)}-v^{(t+1)}),\quad w_{2}^{(t+1)}=w_{2}^{(t+\frac{1}{2})}-s\beta(u^{(t+1)}-\tilde{W}v^{(t+1)}).\end{aligned}\right$. (14)
which can be transformed into the following version ready for decentralized update:
@@
$\left\{\begin{aligned} u^{(t+1)}&=\arg\min_{u\in\mathbb{R}^{nd}}f(u)-\langle\tilde{W}w_{1}^{(t)}+w_{2}^{(t)}+\beta(2\tilde{W}v^{(t)}+(1+\tau)u^{(t)}-\tilde{W}^{\top}\tilde{W}u^{(t)}),u\rangle\\ &+\beta(1+\frac{\tau}{2})\|u\|^{2}\\ w_{1}^{(t+\frac{1}{2})}&=w_{1}^{(t)}-r\beta(\tilde{W}u^{(t+1)}-v^{(t)}),\quad w_{2}^{(t+\frac{1}{2})}=w_{2}^{(t)}-r\beta(u^{(t+1)}-\tilde{W}v^{(t)}),\\ v^{(t+1)}&=\arg\min_{v\in\mathbb{R}^{nd}}g(v)-\langle\beta(2\tilde{W}u^{(t+1)}+(1+\tau)v^{(t)}-\tilde{W}^{\top}\tilde{W}v^{(t)})-(w_{1}^{(t+\frac{1}{2})}+\\ &\tilde{W}w_{2}^{(t+\frac{1}{2})}),v\rangle+\beta(1+\frac{\tau}{2})\|v\|^{2}\\ w_{1}^{(t+1)}&=w_{1}^{(t+\frac{1}{2})}-s\beta(\tilde{W}u^{(t+1)}-v^{(t+1)}),\quad w_{2}^{(t+1)}=w_{2}^{(t+\frac{1}{2})}-s\beta(u^{(t+1)}-\tilde{W}v^{(t+1)}).\end{aligned}\right$. (15)

### v2

@@ 3.3 Optimal Double-Communication Structure
Each iteration of DS-ADMM therefore consists of two local update groups and two neighbor-communication rounds. Information from one group is not used directly in its own update but enables the update of the other group, creating a feedback structure. This interleaving induces a coupled communication-update mechanism where each block drives progress in the other.
@@
Furthermore, this structure makes symmetric ADMM not just suitable but essential: it accelerates convergence without increasing communication, and leads to a clean, symmetric algorithm. To our knowledge, this tightly coupled update-communication framework is novel in the decentralized optimization literature.
@@ 3.4 Algorithm
The resulting procedure is summarized in Algorithm 1 and illustrated in Figure 1, which we refer to as Double-Communication Symmetric ADMM (DS-ADMM). The step-size parameters are set as $0<r\leq 1$ and $s=1$.
Algorithm 1 Double-Communication Symmetric ADMM (DS-ADMM)
1:Initialize: $u_{i}^{(0)}=v_{i}^{(0)}=\lambda_{1,i}^{(0)}=\lambda_{2,i}^{(-\frac{1}{2})}=0$ for all $i\in\{1,\dots,n\}$, mixing matrix $W\in\mathbb{R}^{n\times n}$, step sizes $r>0$, augmented Lagrangian parameter $\beta>0$ and proximal term parameter $\tau>0$.
@@
4:  [Communication 1 ] Transmit $a_{i}^{(t+1)}=\lambda_{2,i}^{(t+\frac{1}{2})}+\frac{1}{r}(\lambda_{2,i}^{(t+\frac{1}{2})}-\lambda_{2,i}^{(t)})$ and $u_{i}^{(t+1)}$.
@@
6:  [Communication 2] Transmit $v_{i}^{(t+1)}$ and $b_{i}^{(t+1)}=2\lambda_{1,i}^{(t+1)}-\lambda_{1,i}^{(t+\frac{1}{2})}$.
@@
Figure 1: Illustration of one iteration of the proposed two-group symmetric decentralized ADMM scheme. Solid arrows denote inter-group communication, i.e., information exchanges that require communication across groups, while dashed arrows indicate intra-group communication performed locally within each group.
@@ 4 Convergence Analysis
which is positive definite. We also define the concatenated variable $\theta=(u^{\top},v^{\top},\lambda^{\top})^{\top}\in\mathbb{R}^{4nd}$.
@@ 4.1 General Sublinear Convergence
Theorems 3.3 and 4.2 of (Gu et al., 2015) imply that the proposed algorithm enjoys a general sublinear convergence rate $\mathcal{O}(1/t)$ without requiring strong assumptions on the objective functions.
@@
Let $\{\theta^{(t)}\}$ be the sequence generated by DS-ADMM. Then $\{\theta^{(t)}\}$ converges to a solution point $\theta^{\infty}$, and the following non-ergodic sublinear rate holds:
$\|\theta^{(t)}-\theta^{(t+1)}\|^{2}\leq\frac{1}{\beta\tau(t+1)}\cdot\frac{1+r}{1-r}\left(\|\theta^{(1)}-\theta^{(0)}\|_{H}^{2}+\|v^{(1)}-v^{(0)}\|_{Q}^{2}\right).$ (7)
@@ 4.2 Linear Convergence under Metric Subregularity
Although various results on linear convergence of ADMM and its variants exist (e.g., (Deng and Yin, 2016; Han et al., 2018; Yuan et al., 2020; Bai et al., 2019; Gu et al., 2015)), none directly apply to our algorithm. Nevertheless, we adapt ideas from these works to establish linear convergence under a standard regularity condition known as metric subregularity.
@@
$T_{\text{KKT}}(\theta):=\begin{pmatrix}\partial f(u)-A^{\top}\lambda\\ \partial g(v)+B^{\top}\lambda\\ Au-Bv\end{pmatrix}$, (9)
and solution set $\Omega^{*}:=\{\theta\mid 0\in T_{\text{KKT}}(\theta)\}$.
@@
Suppose $T_{\mathrm{KKT}}$ is metrically subregular at $(\bar{\theta},0)$ with modulus $c$ for any $\bar{\theta}\in\Omega^{*}$. Then the sequence $\{\theta^{(t)}\}$ generated by DS-ADMM converges Q-linearly to $\Omega^{*}$, i.e., there exist integer $T>0$ and constant $\epsilon>0$ such that for all $t>T$:
$\mathrm{dist}^{2}_{H}(\theta^{(t+1)},\Omega^{*})\leq\frac{1}{1+\epsilon}\cdot\mathrm{dist}^{2}_{H}(\theta^{(t)},\Omega^{*}),$ (10)
@@
$\epsilon=\frac{\phi}{c^{2}\delta\zeta}>0,\quad\phi=\min\left\{2\beta\rho,\frac{1-r}{\beta}\right\},$ (11)
@@
$\delta=\max\left\{6r^{2}+\frac{2}{\beta^{2}},\,12\beta^{2}+4+(\tau\beta)^{2},\,3(\tau\beta)^{2}\right\}\quad\zeta=\frac{2r^{2}\beta^{2}+1}{\beta(r+1)}+(2+\tau-r)\beta.$ (12)
@@ 4.3 Sufficient Conditions for Metric Subregularity
The following proposition gives a characterization of the metric subregularity of $T_{\mathrm{KKT}}$. Each case is justified by different results from the literature: Theorem 46 and Theorem 60 of (Yuan et al., 2020) support the first condition, Robinson’s continuity property (Robinson, 1981) establishes the second, and Lemma 4 of (Latafat et al., 2019) together with Theorem 60 of (Yuan et al., 2020) imply the third.
@@
The KKT mapping $T_{\mathrm{KKT}}$ is metrically subregular at $(\bar{\theta},0)$ for any $\bar{\theta}\in\Theta^{*}$ under any of the following conditions: (i) each $f_{i}$ is smooth and strongly convex, and each $g_{i}$ is PLQ; (ii) all $f_{i}$ and $g_{i}$ are PLQ; (iii) all $f_{i}$ and $g_{i}$ are smooth and strongly convex.
@@ 5 Numerical Experiments
We evaluate the proposed DS-ADMM algorithm on two representative decentralized composite optimization tasks: Lasso regression and $\ell_{2}$-regularized SVM classification. All experiments use $n=30$ agents connected via a random graph with edge probability $p=0.5$. The Metropolis–Hastings mixing matrix is constructed following Appendix A, and data are evenly partitioned among agents. The global regularizer is split uniformly as $g_{i}(x)=\tfrac{1}{n}g(x)$. Comparisons are made against four representative methods: Decentralized Proximal ADMM (Wang et al., 2018), PG-EXTRA (Shi et al., 2015a), NIDS  (Li et al., 2019) and ProxMudag (Ye et al., 2023). All experiments are conducted on a machine equipped with an Intel Core i7-1260P CPU and 16GB RAM.
All adaptive parameters of baseline methods are tuned for best performance. For DS-ADMM, we use fixed step sizes $(r,s)=(0.99,1)$ and proximal coefficient $\tau=0.01$. Suboptimality is measured as $F(\bar{u}^{(t)})-F(u^{\star})$, where $u^{\star}$ is a high-accuracy centralized solution.
Lasso Regression.
Using the a9a dataset from the LIBSVM repository (Chang and Lin, 2011) with
$f_{i}(x)=\frac{1}{2m}\|A_{i}x-b_{i}\|^{2},\qquad g_{i}(x)=\frac{\lambda}{n}\|x\|_{1},\quad\lambda=\tfrac{1}{m}.$
SVM Classification.
Using a1a, each agent solves
$f_{i}(x)=\frac{1}{m}\!\sum_{j\in S_{i}}\!\max(0,1-b_{j}a_{j}^{\top}x),\qquad g_{i}(x)=\frac{\lambda}{2n}\|x\|_{2}^{2}.$
Note that ProxMudag is excluded in the SVM classification task as its formulation requires globally coupled nonsmooth terms. Across both tasks, DS-ADMM consistently converges faster and requires significantly fewer communication rounds. Additional results and parameter sensitivity analysis are reported in Appendix D.
Lasso: iterations
Lasso: communication rounds
SVM: iterations
SVM: communication rounds
Figure 2: Comparison of DS-ADMM with baseline methods on Lasso and SVM.
@@ 6 Conclusion
Acknowledgments
This research is partially supported by the National Natural Science Foundation of China (12231017, 72171216) and the National Key R&D Program of China (2022YFA1003803). We gratefully acknowledge the insightful comments and suggestions provided by the anonymous reviewers.
@@ Appendix B Global Form of Update
Then we decompose the Lagrange multiplier $\lambda$, which yields the below equation:
@@
$\left\{\begin{aligned} u^{(t+1)}&=\arg\min_{u\in\mathbb{R}^{nd}}f(u)-\langle\lambda_{1}^{(t)},\widetilde{W}u\rangle-\langle\lambda_{2}^{(t)},u\rangle+\frac{\beta}{2}\|\widetilde{W}u-v^{(t)}\|^{2}+\frac{\beta}{2}\|u-\widetilde{W}v^{(t)}\|^{2}\\ &+\frac{1}{2}\|u-u^{(t)}\|_{Q}^{2},\\ \lambda_{1}^{(t+\frac{1}{2})}&=\lambda_{1}^{(t)}-r\beta(\widetilde{W}u^{(t+1)}-v^{(t)}),\quad\lambda_{2}^{(t+\frac{1}{2})}=\lambda_{2}^{(t)}-r\beta(u^{(t+1)}-\widetilde{W}v^{(t)}),\\ v^{(t+1)}&=\arg\min_{v\in\mathbb{R}^{nd}}g(v)+\langle\lambda_{1}^{(t+\frac{1}{2})},v\rangle+\langle\lambda_{2}^{(t+\frac{1}{2})},\widetilde{W}v\rangle+\frac{\beta}{2}\|\widetilde{W}u^{(t+1)}-v\|^{2}\\ &+\frac{\beta}{2}\|u^{(t+1)}-\widetilde{W}v\|^{2}+\frac{1}{2}\|v-v^{(t)}\|_{Q}^{2},\\ \lambda_{1}^{(t+1)}&=\lambda_{1}^{(t+\frac{1}{2})}-s\beta(\widetilde{W}u^{(t+1)}-v^{(t+1)}),\quad\lambda_{2}^{(t+1)}=\lambda_{2}^{(t+\frac{1}{2})}-s\beta(u^{(t+1)}-\widetilde{W}v^{(t+1)}).\end{aligned}\right$. (16)
And it can be transformed into the following version ready for decentralized update:
@@
$\left\{\begin{aligned} u^{(t+1)}&=\arg\min_{u\in\mathbb{R}^{nd}}f(u)-\langle\widetilde{W}\lambda_{1}^{(t)}+\lambda_{2}^{(t)}+\beta(2\widetilde{W}v^{(t)}+(1+\tau)u^{(t)}-\widetilde{W}^{\top}\widetilde{W}u^{(t)}),u\rangle\\ &+\beta(1+\frac{\tau}{2})\|u\|^{2}\\ \lambda_{1}^{(t+\frac{1}{2})}&=\lambda_{1}^{(t)}-r\beta(\widetilde{W}u^{(t+1)}-v^{(t)}),\quad\lambda_{2}^{(t+\frac{1}{2})}=\lambda_{2}^{(t)}-r\beta(u^{(t+1)}-\widetilde{W}v^{(t)}),\\ v^{(t+1)}&=\arg\min_{v\in\mathbb{R}^{nd}}g(v)-\langle\beta(2\widetilde{W}u^{(t+1)}+(1+\tau)v^{(t)}-\widetilde{W}^{\top}\widetilde{W}v^{(t)})-(\lambda_{1}^{(t+\frac{1}{2})}+\\ &\widetilde{W}\lambda_{2}^{(t+\frac{1}{2})}),v\rangle+\beta(1+\frac{\tau}{2})\|v\|^{2}\\ \lambda_{1}^{(t+1)}&=\lambda_{1}^{(t+\frac{1}{2})}-s\beta(\widetilde{W}u^{(t+1)}-v^{(t+1)}),\quad\lambda_{2}^{(t+1)}=\lambda_{2}^{(t+\frac{1}{2})}-s\beta(u^{(t+1)}-\widetilde{W}v^{(t+1)}).\end{aligned}\right$. (17)
Rearranging the terms and using the facts that $\widetilde{W}v^{(t)}-\widetilde{W}^{\top}\widetilde{W}u^{(t)}=\frac{1}{s\beta}\widetilde{W}(\lambda_{1}^{(t)}-\lambda_{1}^{(t-\frac{1}{2})})$ and $\widetilde{W}u^{(t+1)}-\widetilde{W}^{\top}\widetilde{W}v^{(t)}=-\frac{1}{r\beta}\widetilde{W}(\lambda_{2}^{(t+\frac{1}{2})}-\lambda_{1}^{(t)})$ in the above equation, we get:
$\left\{\begin{aligned} u^{(t+1)}&=\arg\min_{u\in\mathbb{R}^{nd}}f(u)-\langle\widetilde{W}\lambda_{1}^{(t)}+\lambda_{2}^{(t)}+\beta(\widetilde{W}v^{(t)}+(1+\tau)u^{(t)})+\frac{1}{s}\widetilde{W}(\lambda_{1}^{(t)}-\lambda_{1}^{(t-\frac{1}{2})}),u\rangle\\ &+\beta(1+\frac{\tau}{2})\|u\|^{2}\\ \lambda_{1}^{(t+\frac{1}{2})}&=\lambda_{1}^{(t)}-r\beta(\widetilde{W}u^{(t+1)}-v^{(t)}),\quad\lambda_{2}^{(t+\frac{1}{2})}=\lambda_{2}^{(t)}-r\beta(u^{(t+1)}-\widetilde{W}v^{(t)}),\\ v^{(t+1)}&=\arg\min_{v\in\mathbb{R}^{nd}}g(v)-\langle\beta(\widetilde{W}u^{(t+1)}+(1+\tau)v^{(t)})-(\lambda_{1}^{(t+\frac{1}{2})}+\widetilde{W}\lambda_{2}^{(t+\frac{1}{2})}\\ &+\frac{1}{r}\widetilde{W}(\lambda_{2}^{(t+\frac{1}{2})}-\lambda_{1}^{(t)})),v\rangle+\beta(1+\frac{\tau}{2})\|v\|^{2}\\ \lambda_{1}^{(t+1)}&=\lambda_{1}^{(t+\frac{1}{2})}-s\beta(\widetilde{W}u^{(t+1)}-v^{(t+1)}),\quad\lambda_{2}^{(t+1)}=\lambda_{2}^{(t+\frac{1}{2})}-s\beta(u^{(t+1)}-\widetilde{W}v^{(t+1)}).\end{aligned}\right$. (18)
By our previous definition of $a^{(t)}=\lambda_{2}^{(t+\frac{1}{2})}+\frac{1}{r}(\lambda_{2}^{(t+\frac{1}{2})}-\lambda_{2}^{(t)})$ and $b^{(t)}=\lambda_{1}^{(t)}+\frac{1}{s}(\lambda_{1}^{(t)}-\lambda_{1}^{(t-\frac{1}{2})})$, we get
$\left\{\begin{aligned} u^{(t+1)}&=\arg\min_{u\in\mathbb{R}^{nd}}f(u)-\langle\widetilde{W}a^{(t)}+\lambda_{2}^{(t)}+\beta(\widetilde{W}v^{(t)}+(1+\tau)u^{(t)}),u\rangle+\beta(1+\frac{\tau}{2})\|u\|^{2}\\ \lambda_{1}^{(t+\frac{1}{2})}&=\lambda_{1}^{(t)}-r\beta(\widetilde{W}u^{(t+1)}-v^{(t)}),\quad\lambda_{2}^{(t+\frac{1}{2})}=\lambda_{2}^{(t)}-r\beta(u^{(t+1)}-\widetilde{W}v^{(t)}),\\ v^{(t+1)}&=\arg\min_{v\in\mathbb{R}^{nd}}g(v)-\langle\beta(\widetilde{W}u^{(t+1)}+(1+\tau)v^{(t)})-(\lambda_{1}^{(t+\frac{1}{2})}+\widetilde{W}b^{(t)}),v\rangle+\beta(1+\frac{\tau}{2})\|v\|^{2}\\ \lambda_{1}^{(t+1)}&=\lambda_{1}^{(t+\frac{1}{2})}-s\beta(\widetilde{W}u^{(t+1)}-v^{(t+1)}),\quad\lambda_{2}^{(t+1)}=\lambda_{2}^{(t+\frac{1}{2})}-s\beta(u^{(t+1)}-\widetilde{W}v^{(t+1)}).\end{aligned}\right$. (19)